# ManiSkill3: GPU Parallelized Robotics Simulation and Rendering for Generalizable Embodied AI

### v1

@@ 3.3 Heterogeneous GPU Simulation
ManiSkill3 is so far the only simulation framework that supports heterogeneous GPU simulation. This is the feature of being able to simulate different object geometries, different numbers of objects, and different articulations with different DOFs across different parallel environments. For example, in the Open Cabinet Drawer task, for each parallel environment, we build a different cabinet (all with different DOFs) and sample a random drawer link that needs to be opened to succeed. In the Pick Clutter YCB task, we sample a different number of YCB objects in each parallel environment and sample one random object out of the clutter that is meant to be picked up. ManiSkill3 easily supports this kind of simulation and further supports rendering these different scenes in parallel all at once with an example 3rd view rendering, as shown in Figure 6. Heterogeneous GPU simulation enables more generalizable learning for manipulation as algorithms like those in RL can now simultaneously train on every single object from the YCB object dataset (Çalli et al., 2015) or the PartNet Mobility Dataset of cabinets (Mo et al., 2019).
3.4 Sim2Real and Real2Sim for Robot Manipulation
@@
Figure 7: Three different kinds of digital twins in ManiSkill3. Top row shows the real-world setup and bottom row shows the digital twin. Left: Digital twin of a cube picking task. Middle: Digital twin of the vision-tactile simulation of a key insertion task. Right: Real2sim digital twin of spoon placing task.
@@
Towards the goal of robust real-world robotics beyond simulation, we verify sim2real and real2sim are both possible using ManiSkill3 via digital twins on some tasks. Figure 7 showcases several digital twins supported with real world counterparts. For the cube picking task, we segment out relevant objects to feed into a PointNet model and train with PPO on only simulated data, ultimately achieving a real world success rate of 95% over a 50cm x 50cm workspace. For the vision-tactile peg insertion task, we simulate the tactile sensor made of silicone as a softbody and refer readers to the results showcased in the original work by Chen et al. (2024) for those environments, which achieved a 95.08% success rate in the real world. Finally, for the real2sim digital twins we evaluate Octo and RT-1X on the ManiSkill3 GPU parallelized version of 4 tasks in SIMPLER (Li et al., 2024). We achieve a correlation between real-world success rates and simulation success rates of 0.9284 and a Mean Maximum Rank Violation (MMRV) value 0.0147, which is close to the original values reported in SIMPLER. See Appendix A.11 for more details on real2sim evaluations.
3.5 Simple Unified API for Building GPU Simulated Robotics Tasks
@@
A core reason behind the flexibility of the ManiSkill3 system to support so many different distinct task categories is the clean and simple API for task-building. The simple API also enables users to easily build and customize their own robotics tasks.
3.5.1 Object-Oriented API for Articulations, Links, Joints, and Actors
@@ 3.5.2 Robots and Controllers
ManiSkill3 supports both URDF and Mujoco MJCF definition formats natively and builds articulated robots based on the URDF/MJCF directly. For each robot, ManiSkill3 further provides a number of configurable controller options for both GPU parallelized joint position control and inverse-kinematic (IK) control, modified from ManiSkill2 for GPU simulation. ManiSkill3 builds upon the PyTorch Kinematics package (Zhong et al., 2024) to support inverse-kinematic based controllers parallelized on the GPU. These options are easily configured at runtime with either preset configurations or user-supplied controller configurations. Currently, there are 20+ different robots supported out of the box in ManiSkill3, a subset of which are visualized in Figure 11 in the Appendix. Finally, ManiSkill3 comes with extensive tutorials and examples of how to tune and optimize robots for fast simulation, which has often proven a stumbling block for those new to robot simulation importing complex robots for the first time. The tutorials highlight details from how to simulate mobile-bases to simulation hyperparameter tuning for improving simulation speed and/or accuracy.
3.6 Demonstration Datasets
@@
We leverage a variety of approaches to collect/generate our demonstration datasets infinitely at scale. For the simplest tasks, we write and open source some motion planning-based solutions to generate demonstration data. Some tasks with easy-to-define reward functions have dense reward functions defined and converged RL policies are used to generate demonstration data. For more difficult tasks, we collect demonstration data (typically about 10 demonstrations) via teleoperation tools. Then, we use RFCL (Tao et al., 2024) or RLPD (Ball et al., 2023) to run fast online imitation learning and generate data from converged policies.
We further adapt the trajectory replay tool from ManiSkill2 to work with both CPU and GPU simulated demonstration data. The replay tool enables users to change the observations stored (e.g., state or rgbd), as well as modify the actions stored (e.g. joint position or delta end effector pose). The tool makes working with demonstration data more flexible in ManiSkill3, such as converting joint position actions generated by motion planning to other control modes like end-effector control. We are currently still in the process of generating demonstrations for more tasks.
4 Baselines and Results
ManiSkill3 provides several popular robot learning baselines that span 4 different categories:
@@ 4 Baselines and Results
Wall-time Efficient Reinforcement Learning: We include a torch based vectorized implementation of model-free RL algorithms PPO and SAC (Haarnoja et al., 2018), as well as the state-of-the-art model-based RL algorithm TD-MPC2 (Hansen et al., 2024). Configurations for baselines are tuned to minimize training wall time with no regard to sample efficiency.
@@
Sample Efficient Reinforcement Learning: All of the RL baselines in the wall-time efficient setting are included here with configurations tuned towards more gradient updates and fewer environment steps to maximize sample efficiency.
Offline Imitation Learning: These are supervised learning and offline RL style baselines. We currently provide Behavior Cloning, offline variants of SAC, and the current state-of-the-art Diffusion Policy (Chi et al., 2023) as baselines.
@@
Online Imitation Learning: Online imitation learning generally refers to algorithms that learn from demonstrations in addition to collecting online environment transitions. We currently provide the two state-of-the-art baselines in this area, Reinforcement Learning from Prior Data (RLPD) (Ball et al., 2023), and Reverse Forward Curriculum Learning (RFCL) (Tao et al., 2024). Note that RFCL leverages simulation state resets.
Furthermore, ManiSkill3 provides tools that establish clear specifications for how to evaluate policies that address common issues in other simulators that can often cause confusion.
4.1 Reinforcement Learning Training/Evaluation Setup
As part of an effort to unify baselines and robot simulation to allow easy comparison and research/testing, we ensure that all baselines report the same metrics and run the same evaluation environment setups. With shared metrics, this enables us to form a massive Weights and Biases (Wandb) online dashboard of every RL (and IL) algorithm’s results and allows for easy cross-comparison.
We observe that in past simulators/research, there is often inconsistency in how episode truncations/terminations are handled and how to report environment success/fail. For example, Isaac Lab by default terminates episodes early during training before reaching the timelimit if the agent succeeds or fails, but many RL algorithms (especially off-policy ones) like TD-MPC2 and SAC model environments as having infinite horizon and do not terminate environments early due to success/fail. Moreover, Isaac Lab is inconsistent with whether termination means success or failure in different environments, as shown in their Franka Open Cabinet code and Anymal C Locomotion code.
To ensure better consistency in metrics reported not only by baselines provided in ManiSkill3, but also those reported by users, we provide an environment wrapper for evaluation environments that auto-records these defined metrics. We provide clear instructions on how to set it up, record results, and exactly what each metric means. Concretely, the wrapper ensures evaluation of policies never terminate environments early (it is permitted during training as early termination can accelerate some RL algorithms), and the wrapper records episode return, length, success/fail once, and success/fail at end metrics.
Success once is equivalent to early termination due to success and is set true if the environment achieves success once at any point during an episode before truncation. Success at the end is equivalent to recording whether the environment is in a success state at the final episode step before truncation. Similarly, fail once/at the end operate the same way but for failure conditions. Both “once” and “at the end” metrics are recorded to indicate a clear distinction in the types of success/failure a policy might achieve to avoid confusion.
Finally, we make a key distinction in sample-efficient and wall-time-efficient training results by explicitly tagging which training runs are using sample/wall-time-efficient configurations. Research into both sample efficiency and wall-time efficiency remains active in the robot learning community so we ensure this distinction is clearly labelled. The final results on the Wandb dashboard are still being uploaded.
4.2 Imitation Learning Training/Evaluation Setup
@@
Mandlekar et al. (2021) show that imitation learning algorithm performance heavily depends on how the demonstrations were collected, particularly on how “multimodal” the data is. For example, many behavior cloning algorithms regress actions and perform poorly when trained on motion planning or human teleoperated data, but can perform very well if trained on data generated by a deterministic neural network. With this caveat in mind, we explicitly track in all our imitation learning (online and offline) baselines how many demonstrations are used (an integer), what type of demonstrations are used (neural net, motion planning, or human), and where the demonstrations are sourced from exactly (a longer description e.g. neural net trained via TD-MPC2, teleoperation via Meta-Quest VR). On the ManiSkill3 Weights and Biases online dashboard, these key configurations over the dataset can be searched/filtered, enabling users to easily compare algorithm performance on any demonstration dataset ablation they wish.
Furthermore, we provide shared code/evaluation harnesses that enable users to easily use a vectorized CPU backend or the GPU simulation backend for fast evaluation of imitation learning policies during training. Success/fail once/at the end metrics are similarly also required and reported automatically following the same evaluation setup as RL algorithms. The final results on the Wandb dashboard are still being uploaded.
4.3 Training Speed
@@
We run experiments using the Proximal Policy Optimization (PPO) (Schulman et al., 2017) RL algorithm on the ManiSkill3 GPU simulation and the ManiSkill2 CPU simulation. ManiSkill2 was previously the fastest robotics simulation+rendering framework until ManiSkill3. The experiments were run on an RTX-4090 GPU on the PickCube task, where a Franka robot arm must grasp a randomly initialized cube and hold it still at a random goal location. For the vision-based experiment no ground truth data like cube pose is provided and the RL policy must solve from proprioceptive information and a single 128x128 resolution RGB image rendered by the environment’s 3rd-person camera. RL hyperparameters are tuned to achieve the fastest training time in both settings. Results in Figure 10 show that state and vision based training are massively accelerated with GPU simulation and rendering.
We note that while the speedups are around 5x in the results, if more scalable RL algorithms are used the speedups can be greater as PPO has worse sample-efficiency when using a high number of parallel environments compared to typical CPU simulation setups with less parallel environments.
Figure 10: Training time and sample-efficiency of PPO on GPU/CPU simulation.
5 Conclusion
@@
ManiSkill3 introduces a state-of-the-art framework/benchmark for generalizable robotics simulation and rendering. ManiSkill3 runs faster, uses less GPU memory, and supports the most diverse range of robotics tasks compared to alternative simulators. In particular, ManiSkill3 offers unparalleled support for both sim2real and real2sim environments in manipulation tasks. We believe that this comprehensive approach will encourage the research community to tackle manipulation challenges more extensively through compute-scalable simulation. Furthermore, ManiSkill3 provides an easy-to-use object-oriented API for building GPU simulated heterogeneous tasks, democratizing access to scalable robot learning. Finally, demonstrations and RL/IL baselines with clearly defined metrics are open sourced for users to use.

### v2

@@ III-C Heterogeneous GPU Simulation
ManiSkill3 is so far the only simulation framework that completely supports heterogeneous GPU simulation. This is the feature of being able to simulate different object geometries, different numbers of objects, and different articulations with different DOFs across parallel environments. For example, in the OpenCabinetDrawer task, for each parallel environment, we build a different cabinet (all with different DOFs) and sample a random drawer link that needs to be opened to succeed. In the Pick Clutter YCB task, we sample a different number of YCB objects in each parallel environment and sample one random object out of the clutter as the goal object to pick. ManiSkill3 easily supports this kind of simulation and further supports rendering these different scenes in parallel all at once with an example 3rd view rendering illustrated in Figure 6. Heterogeneous GPU simulation enables more generalizable learning for manipulation as algorithms like PPO can simultaneously train on every single object from the YCB dataset [5] or the PartNetMobility dataset of cabinets [37].
III-D Teleoperation
We provide a Virtual Reality (VR) Teleoperation system that is seamlessly integrated into ManiSkill3 with immersive visual feedback and supports wired connections. ManiSkill3 receives real-time hand pose data from the teleoperator that is translated into corresponding robot actions. At the same time, the system streams 4K stereo video via Air Light VR (ALVR) to the VR device at 60 Hz, ensuring a smooth and immersive ”all-scene view” experience. This allows the user to explore the entire environment freely, making it possible to teleoperate long-horizon tasks as well as more precise tasks by simply looking closer in the simulation. A crucial control feature is “where your hand is, where the end-effector is”, which tightly aligns the operator’s hand movement with the robot’s end-effector. Prior work provide ways to stream video feedback over the internet via a hardware agnostic web app [11, 15] but tradeoff some ease-of-use for better accessibility (such as long-range remote teleoperation). Our integrated system enables wider viewports and reduce the need for teleoperators to compensate for hand-robot misalignment and visual disparities, visualized in 7. Our VR teleop system is also compatible with real-world teleoperation. For real-world operations, we use one or more depth cameras to generate a point cloud of the scene. The point cloud is then rendered and streamed as stereo RGB video to the operator in real time, allowing for effective and immersive control. Example code for configuring this system is provided for teleoperating robot arms or floating robots with multiple grippers/fingers. See Appendix XII for more details on this teleoperation system.
Figure 7: Visualization of VR teleoperation system. Left: A teleoperator using hand poses captured by the Meta Quest 3 headset to control robot motion in real-time. Middle: A 360-degree scene displayed in the VR device, providing immersive sensory feedback. Right: Trajectory replay.
III-E Sim2Real and Real2Sim for Robot Manipulation
@@
Figure 8: Three different kinds of digital twins in ManiSkill3. Top row shows the real-world setup and bottom row shows the digital twin. Left: Domain randomized digital twin of a cube picking task. Middle: Digital twin of the vision-tactile simulation of a key insertion task. Right: Real2sim digital twin of a spoon placing task.
@@
Towards the goal of robust real-world robotics beyond simulation, we verify sim2real and real2sim are both possible using ManiSkill3 via digital twins on some tasks. Figure 8 showcases several digital twins supported with real world counterparts. For the cube picking task, we train with RL on simulation images and directly deploy to the real world, achieving a real world success rate of 91.6% averaged across 3 RL training runs. See Section 12 for more details on rigid body manipulation sim2real. For the vision-tactile peg insertion task, we simulate the tactile sensor made of silicone as a softbody and refer readers to the results showcased in the original work by [9] for those environments, which achieved a 95.08% success rate in the real world. Finally, for the real2sim digital twins we evaluate Octo and RT-1X on the ManiSkill3 GPU parallelized version of 4 tasks in SIMPLER [31]. We achieve a correlation between real-world success rates and simulation success rates of 0.9284 and a Mean Maximum Rank Violation (MMRV) value 0.0147, which is close to the original values reported in SIMPLER. See Appendix VII-K for more details on real2sim evaluations.
III-F Simple Unified API for Building GPU Parallelized Robotics Tasks
@@
A core reason behind the flexibility of the ManiSkill3 system to support so many different distinct task categories is the clean and simple API for task-building. The simple API also enables users to easily build and customize their own robotics tasks without having to worry about complex GPU memory management details or desigining robot controllers. In addition to the API design extensive tutorials are provided for customizing tasks. We describe two general features provided by ManiSkill3 below:
III-F1 Object-Oriented API for Articulations, Links, Joints, and Actors
@@ III-F2 Robots and Controllers
ManiSkill3 supports both URDF and Mujoco MJCF definition formats natively and builds articulated robots based on the URDF/MJCF directly. For each robot, ManiSkill3 further provides a number of pre-built configurable controller options for both GPU parallelized joint position control and inverse-kinematic (IK) control, modified from ManiSkill2 for GPU simulation. ManiSkill3 builds upon the PyTorch Kinematics package [55] to support inverse-kinematic based controllers parallelized on the GPU, typically used to control robot arm end-effectors. These options are easily configured at runtime with either preset configurations or user-supplied controller configurations. Currently, there are 20+ different robots supported out of the box in ManiSkill3, a subset of which are visualized in Figure 14 in the Appendix. Finally, ManiSkill3 comes with extensive tutorials and examples of how to tune and optimize robots for fast simulation, which has often proven a stumbling block for those new to robot simulation importing complex robots for the first time.
III-G Demonstration Datasets
@@
We leverage a variety of approaches to collect/generate our demonstration datasets infinitely at scale. For the simplest tasks, we write and open source some motion planning-based solutions to generate demonstration data. Some tasks with easy-to-define reward functions have dense reward functions defined and converged RL policies are used to generate demonstration data. For more difficult tasks, we collect demonstration data (typically about 10 demonstrations) via teleoperation tools. Then, we use RFCL [47] or RLPD [2] to run fast online imitation learning and generate data from converged policies.
We further adapt the trajectory replay tool from ManiSkill2 to work with both CPU and GPU simulated demonstration data. The replay tool enables users to change the observations stored (e.g., state or rgbd, and allow modifying the rendering shaders used) as well as modifying the rewards stored (dense or sparse). In tasks involving the Franka robot arms, we provide code to convert actions from one controller type to another (e.g. joint position control to delta end effector pose control). Importantly we support collecting data in CPU/GPU simulation and replaying them in CPU/GPU simulation with different numbers of parallel environments via explicit control over randomization/RNG seeding at the per-parallel-environment level. This enables flexibility in trajectory replay as data collected on one machine with more GPU memory can be replayed on other machines with less GPU memory that cannot use as many parallel environments.
IV Baselines and Results
ManiSkill3 provides several popular robot learning baselines as well as simple reproducible setups for end-to-end trainable vision-based sim2real policies.
IV-A Reinforcement Learning
We provide two categories of RL baselines as follows:
@@ IV-A Reinforcement Learning
Wall-time Efficient Reinforcement Learning: We include a torch based vectorized implementation of model-free RL algorithms PPO and SAC [20], as well as the state-of-the-art model-based RL algorithm TD-MPC2 [22]. Configurations for baselines are tuned to minimize training wall time with no regard to sample efficiency. Code for PPO and SAC is implemented based on CleanRL [23] and leverages the torch compile and cudagraphs acceleration features introduced by LeanRL.
@@
Sample Efficient Reinforcement Learning: All of the RL baselines in the wall-time efficient setting besides PPO are included here with configurations tuned towards more gradient updates and fewer environment steps to maximize sample efficiency.
We share results on training speed in Section IV-C. More in-depth details on RL setups and more training results are shared in Appendix IX.
IV-B Learning From Demonstrations (LfD) / Imitation Learning
We provide two categories of LfD baselines as follows:
Offline Imitation Learning: We currently provide Behavior Cloning, Diffusion Policy [13], Action Chunking Transformer [54], and PerACT [45] as baselines. We also support evaluating (but not training) several vision-language action (VLA) models, namely Octo [40], RT-X [14], and RDT-1B [32]. We leave to future work to support training VLA models on simulation data as RDT-1B has done with the previous ManiSkill2 datasets.
@@ IV-B Learning From Demonstrations (LfD) / Imitation Learning
Online Imitation Learning: Online imitation learning generally refers to algorithms that learn from demonstrations in addition to collecting online environment transitions. We currently provide the two state-of-the-art baselines: Reinforcement Learning from Prior Data (RLPD) [2], and Reverse Forward Curriculum Learning (RFCL) [47]. Note that RFCL leverages simulation state resets.
@@
Mandlekar et al. [34] show that imitation learning algorithm performance heavily depends on how the demonstrations were collected, particularly on how “multimodal” the data is. For example, some behavior cloning algorithms perform poorly when trained on motion planning or human teleoperated data, but can perform very well if trained on data generated by a deterministic neural network. With this caveat in mind, we explicitly track in all our LfD baselines the number of demonstrations used (an integer), what type of demonstrations are used (RL generated, motion planning, or human), and where the demonstrations are sourced from exactly (a longer description e.g. neural net trained via TD-MPC2, teleoperation via Meta-Quest VR). Imitation learning results on some environments are shared in Appendix X.
IV-C RL Training Speed
@@
We run experiments using PPO [44] on the ManiSkill3 GPU simulation and the ManiSkill2 CPU simulation. ManiSkill2 was previously the fastest robotics simulation+rendering framework until ManiSkill3. The experiments were run on an RTX-4090 GPU on the PickCube task, where a Franka robot arm must grasp a randomly initialized cube and hold it still at a random goal location. For the vision-based task no ground truth data like cube pose is provided and the RL policy must solve from proprioceptive information and one 128x128 RGB image rendered by the environment’s 3rd-person camera. RL hyperparameters are tuned to achieve the fastest training time in both settings. Results in Figure 11 show that state and vision based training are massively accelerated with GPU simulation and rendering. PickCube with delta joint position control from state-based observations in GPU simulation reaches near 100% success rate after about 1 minute of training, a 15x speed up relative to ManiSkill2. From RGB observations with parallel rendering PickCube is solved after about 10 minutes of training, a 8x speed up relative to ManiSkill2. For results and details of RL on more environments see Appendix IX.
Figure 11: Wall-clock training time of PPO on GPU/CPU simulation showing the average success rate over time across 5 seeds. Shaded areas correspond to the 95% confidence interval.
IV-D Vision-Based Sim2Real Manipulation
Figure 12: Green screening visualized with Koch-v1.1 robot. Left: Real RGB observations. Right: Sim RGB observations overlaid on real background. Middle: Sim RGB observations overlaid on real RGB observations for alignment visualization.
To showcase the use-case of heterogeneous simulation and parallel rendering, we provide a simple reproducible setup for end-to-end training a vision-based manipulation policy that deploys successfully zero-shot. The setup uses the low-cost \$300 Koch robot arm and a phone camera for third-person RGB observations, leveraging the open-source Hugging Face LeRobot library [4] for the hardware and setup to enhance reproducibility and accessibility. This setup is not limited to the Koch arm; other more expensive robotic arms can be used and likely perform better due to more precise hardware.
During simulation training and real-world evaluation, observations are restricted to RGB inputs and robot joint positions; no demonstrations or privileged state information such as cube pose is used, and the robot is controlled via continuous joint level input. We note however that future work combining simulation training with teleoperated demonstrations could improve performance and training time. To match the real background within simulation and reduce the visual sim2real gap we green screen the real-world background. We skip green-screening over the dynamic objects like the cube and robot arm via the environment’s segmentation map, see Figure 12 for an example. To enhance generalization and accommodate variations in real-world setups, we domain randomize a number of elements of the simulation. Task agnostic domain randomizations include camera pose, lighting direction, and the robot pose. Specific to the cube picking task we further domain randomize the cube size, color, and friction.
In this demonstration we picked a random table in a house and set up the green-screening and roughly aligned the simulation camera pose with the real-world camera pose. Then we train on the cube picking environment with PPO using a simple Nature CNN backbone for image feature processing for 15 million samples using 256 parallel environments, taking approximately 1 hour on a RTX 4090 GPU. The trained policy is then zero-shot deployed on the real robot, using the same controller the policy was trained on in simulation, namely a delta target joint position controller. Evaluating the final checkpoint across 3 training runs in the real-world 8 times each yielded a $22/24=91.6\%$ success rate. Real world evaluations test on cubes of varying sizes, colors, and start poses, demonstrating our vision-based sim2real setup is capable of learning robust manipulation policies. Evaluation videos can be found on our demo gallery page.
We further evaluated each intermediate checkpoint from the 3 RL training runs and plot the simulation and real-world success rates in Figure 13. The figure shows a good correlation between real-world and simulated success rates, indicating that some sim2real digital twins built with ManiSkill can be fairly accurate and reflect the real-world. Interestingly, we observe that despite not modelling shadows or the exposed wiring of the robot arm accurately, a successful sim2real policy is still capable of being trained. See Appendix VIII for more details on the exact setup, domain randomizations, controller implementation, and more.
Figure 13: Koch pick-cube sim and real success rates on the grasp cube subtask as well as the full success consisting of grasping, lifting, and return the cube to a goal position. 14 training checkpoints for each of the 3 seeds are evaluated on both sim and real 8 times, using a variety of cube colors and sizes. Success rates are averaged across trials and shaded areas represent 95% confidence intervals.
V Limitations
While parallel rendering can enable some new and more accessible approaches to sim2real and visual RL, it has its limits. For complex environments with many geometries (e.g. room-scale scenes) we can GPU parallel render and simulate them but at a lower number of environments on one GPU.
While it is now possible to do some more generalizable zero-shot RGB based sim2real with a relatively simple setup, some reward engineering was required to encourage safer and more robust grasp behaviors of RL trained policy. Moreover, this is far from solving sim2real completely, but more of opening a new avenue of potential approaches based on fast visual RL in simulation. The sim2real demo is also limited to using just static cameras and it can be interesting future work to explore how to use methods like Gaussian Splatting for novel-view rendering to support mounted cameras. One could also explore algorithnmic changes and/or domain randomizations that can help remove the need for green screening.
Not all types of environments in ManiSkill are readily ”GPU parallelized” in the sense that there are a batch of parallel environments. Most GPU parallelized environments are rigid-body based. The soft body environments are not batched as they use a significant portion of the GPU to simulate just a single environment fast. The vision-tactile simulator that has been tuned heavily for sim2real transfer uses a different set of algorithms for physics simulation compared to the majority of rigid-body only environments.
VI Conclusion
@@
ManiSkill3 introduces a state-of-the-art framework/benchmark for generalizable robotics simulation and rendering. ManiSkill3 uses less GPU memory and depending on scenario can run faster while also supporting heterogeneous GPU simulation. Additionally we support the most diverse range of robotics tasks compared to alternative simulators. Importantly, we reliably support both sim2real and real2sim environments in manipulation tasks with real-world reproducible results. Moreover, we provide a immersive VR teleoperation system option for users to collect data for their own research. Furthermore, ManiSkill3 provides an easy-to-use object-oriented API for building all kinds of GPU simulated tasks, democratizing access to scalable robot learning. Finally, demonstrations and RL/IL baselines with clearly defined metrics are open sourced for users to use. We believe that our comprehensive approach to building the open-osurce ManiSkill3 will encourage the research community to tackle manipulation challenges more extensively through compute-scalable simulation.